# Moments and the Basic Equilibrium Equations

### v1

@@ Appendix A Constraints for Mutual Distances
The HLK theorem and corollary 2.6 play a relevant role in obtaining constraints for mutual distances of configurations. We separate these constraints into two sets: the direct (and simpler) ones that follow directly from the HLK theorem, and the constraints obtained by taking determinants of special matrices associated with a given configuration.
@@ A.1. Direct constraints
Corollary A.1.1 allows us to generalize our early remarks on the rhombus in many ways. For instance, it implies that the difference between the sums of the squared distances from any point to the vertices of two concentric regular $n$-gons of respective squared radii $s$ and $s^{\prime}$ is always equal to $n(s-s^{\prime})$, that is, it depends neither on the point nor on the relative position of the $n$-gons.
@@ A.2. Determinantal constraints
where $\bm{1}$ is the $n\times 1$ matrix $[1\cdots 1]^{T}$. We see from equation (A.2.1) that $W\in\mathbb{W}_{0}(X)$ if and only if $\widetilde{W}=(w_{0},w_{1},\ldots,w_{n})\in\text{Ker}\,\,\mathcal{C}(X)$, with $w_{0}=-\sum_{j=1}^{n}w_{j}\overrightarrow{x_{i}x_{j}}^{2}$, for any $i$. The determinant of $\mathcal{C}(X)$ is known as the Cayley-Menger determinant of $X$ [8, 9].
@@
Theorem A.2.1 implies that $\mathcal{C}(X)$ has nullity equal to $\dim\mathbb{W}_{0}(X)$, which is $c$, by lemma 4.1. For $c=0$, the rank of $\mathcal{C}(X)$ is thus $n+1$, so $\det\mathcal{C}(X)\neq 0$. This means that there are no constraints on mutual distances, as expected. For configurations of codimension $c>0$, $\mathcal{C}(X)$ has rank $n-c+1$, so all subdeterminants of $\det\mathcal{C}(X)$ of size $(n-c+2)\times(n-c+2)$ must vanish. In particular, the $(n-c+2)\times(n-c+2)$ principal minors of $\mathcal{C}(X)$, which are precisely the Cayley-Menger determinants of the Dziobek subconfigurations of $X$ considered in section 4, are all equal to zero. We shall soon check that not all Cayley-Menger determinants of Dziobek subconfigurations are independent and that, in general, a much smaller number of such determinants suffices to guarantee the correct codimension for $X$.
Let us compute the number of constraints that must be satisfied by the mutual distances of a configuration of $n$ points and dimension $d$. Let $E(d)$ be the group of Euclidean (rigid) motions in $\mathbb{R}^{d}$. The manifold of congruence classes of configurations of $n$ points in $\mathbb{R}^{d}$ modulo isometries is $\mathbb{R}^{dn}\big{/}E(d)$, which has dimension $dn-d(d+1)/2$. The natural coordinates to describe the congruence classes are the ($n\choose 2$ mutual distances between the points. For configurations of codimension $c=(n-1)-d>0$, the number of mutual distances exceeds the dimension of the manifold of congruence classes by
(A.2.4) (${n\choose 2}-\left[dn-\frac{d(d+1)}{2}\right]={c+1\choose 2},$
@@
Let $X$ be a configuration of dimension $d$. The mutual distances of $X$ are such that (${c+1\choose 2}$ functionally independent Cayley-Menger determinants of subconfigurations of $X$ vanish, where $c=\emph{codim}(X)$.
@@
Inspired by the argument in section 3.5 of [21], we use induction on the codimension of $X$. When $c=1$, there is the unique constraint $\det\mathcal{C}(X)=0$, so the formula is valid in this case. Suppose the formula holds for a configuration $X$ of $n$ points and dimension $d$, where $d=n-(c+1)$. Adding the point $p$ to $X$ without changing the dimension, the new configuration will have codimension $n+1-(d+1)=n-d=c+1$. The additional constraints are the Cayley-Menger determinants of subconfigurations of $X\cup\{p\}$ formed by a selection of (any) $d$ points in $X$, the point $p$, and each of the remaining $n-d$ points of $X$, one at a time. Thus we get a total of
(${c+1\choose 2}+n-d={c+2\choose 2}$
@@ Appendix B Extended Leibniz Identities
In this appendix we show that the kernel of $\mathcal{B}(X)$, viewed as a symmetric bilinear form on the space $\mathbb{D}=\mu_{0}^{-1}(0)$ of weight vectors with $\mu_{0}=0$, is precisely $\mathbb{W}_{0}(X)$, the space of weight vectors with $\mu_{1}=\overrightarrow{O}$. This basic remark, which appears in [3], produces relations of interest to applications. Here we provide an independent proof, followed by an application to central configurations.
@@
where $\bm{0}$ is the $n\times 1$ matrix $[0\ldots 0]^{T}$. Let $W_{2}\in\mathbb{D}$. Multiplying both sides of the latter equation on the left by $W_{2}^{T}$ and using that $W_{2}^{T}\bm{1}=\bm{1}^{T}W_{2}=0$, we arrive at the extended Leibniz identity

### v2

@@ Appendix A Constraints for Mutual Distances
The HLK theorem and corollary 2.6 play a relevant role in obtaining constraints for mutual distances of configurations. We separate these constraints into two sets: the direct (and simpler) ones that follow straight from the HLK theorem, and the constraints obtained by taking determinants of special matrices associated with a given configuration.
@@ A.1. Direct constraints
Corollary A.1.1 allows us to generalize our earlier remarks on the rhombus in many ways. For instance, it implies that the difference between the sums of the squared distances from any point to the vertices of two concentric regular $n$-gons of respective squared radii $s$ and $s^{\prime}$ is always equal to $n(s-s^{\prime})$, that is, it depends neither on the point nor on the relative position of the $n$-gons.
@@ A.2. Determinantal constraints
where $\bm{1}$ is the $n\times 1$ matrix $1\cdots 1]^{T}$. We see from equation (A.2.1) that $W\in\mathbb{W}_{0}(X)$ if and only if $\widetilde{W}=(w_{0},w_{1},\ldots,w_{n})\in\text{Ker}\,\,\mathcal{C}(X)$, with $w_{0}=-\sum_{j=1}^{n}w_{j}\overrightarrow{x_{i}x_{j}}^{2}$, for any $i$. The determinant of $\mathcal{C}(X)$ is known as the Cayley-Menger determinant of $X$ [8, 9].
@@
Theorem A.2.1 implies that $\mathcal{C}(X)$ has nullity equal to $\dim\mathbb{W}_{0}(X)$, which is $c$, by lemma 4.1. For $c=0$, the rank of $\mathcal{C}(X)$ is thus $n+1$, so $\det\mathcal{C}(X)\neq 0$. This means that there are no constraints on mutual distances, as expected. For configurations of codimension $c>0$, $\mathcal{C}(X)$ has rank nc+1n-c+1, so all subdeterminants of $\det\mathcal{C}(X)$ of size $(n-c+2)\times(n-c+2)$ must vanish. In particular, the $(n-c+2)\times(n-c+2)$ principal minors of $\mathcal{C}(X)$, which are precisely the Cayley-Menger determinants of the Dziobek subconfigurations of $X$ considered in section 4, are all equal to zero. We shall soon check that not all Cayley-Menger determinants of Dziobek subconfigurations are independent and that, in general, a much smaller number of such determinants suffices to guarantee the correct codimension for $X$.
Let us compute the number of constraints that must be satisfied by the mutual distances of a configuration of $n$ points and dimension $d$. Let $E(d)$ be the group of Euclidean (rigid) motions in $\mathbb{R}^{d}$. The manifold of congruence classes of configurations of $n$ points in $\mathbb{R}^{d}$ modulo isometries is $\mathbb{R}^{dn}\big{/}E(d)$, which has dimension dnd(d+1)/2dn-d(d+1)/2. The natural coordinates to describe the congruence classes are the $n\choose 2$ mutual distances between the points. For configurations of codimension c=(n1)d>0c=(n-1)-d>0, the number of mutual distances exceeds the dimension of the manifold of congruence classes by
(A.2.4) ${n\choose 2}-\left[dn-\frac{d(d+1)}{2}\right]={c+1\choose 2}$,
@@
Let $X$ be a configuration of dimension $d$. The mutual distances of $X$ are such that $c+1\choose 2$} functionally independent Cayley-Menger determinants of subconfigurations of $X$ vanish, where $c=\emph{codim}(X)$.
@@
Inspired by the argument in section 3.5 of [21], we use induction on the codimension of $X$. When $c=1$, there is the unique constraint $\det\mathcal{C}(X)=0$, so the formula is valid in this case. Suppose the formula holds for a configuration $X$ of $n$ points and dimension $d$, where d=n(c+1)d=n-(c+1). Adding the point $p$ to $X$ without changing the dimension, the new configuration will have codimension n+1(d+1)=nd=c+1n+1-(d+1)=n-d=c+1. The additional constraints are the Cayley-Menger determinants of subconfigurations of $X\cup\{p\}$ formed by a selection of (any) $d$ points in $X$, the point $p$, and each of the remaining ndn-d points of $X$, one at a time. Thus we get a total of
${c+1\choose 2}+n-d={c+2\choose 2}$
@@ Appendix B Extended Leibniz Identities
In this appendix we show that the kernel of $\mathcal{B}(X)$, viewed as a symmetric bilinear form on the space $\mathbb{D}=\mu_{0}^{-1}(0)$ is precisely $\mathbb{W}_{0}(X)$, the space of weight vectors with $\mu_{1}=\overrightarrow{O}$. This basic remark, which appears in [3], produces relations of interest to applications. Here we provide an independent proof, followed by an application to central configurations.
@@
where $\bm{0}$ is the $n\times 1$ matrix $0\ldots 0]^{T}$. Let $W_{2}\in\mathbb{D}$. Multiplying both sides of the latter equation on the left by $W_{2}^{T}$ and using that $W_{2}^{T}\bm{1}=\bm{1}^{T}W_{2}=0$, we arrive at the extended Leibniz identity